OpenStudy (anonymous):

x^2 - 9x +8 = 0 solving quadratic equations by factoring

OpenStudy (tkhunny):

Really Big Hint: (-1)(-8) = 8

OpenStudy (anonymous):

You need to find two numbers which when added together will give you b (i.e. -9) and when multiplied together will give you c (i.e. 8) recall: x^2 -9x + 8 is of the form ax^2 + bx + c and in your case: a=1, b=-9, c=8 and as tkhunny says: the numbers (-1) and (-8) will work since if you add those together you will get -9 and if you multiply those together you will get 8. Therefore your new equation will look like (x-1)(x-8)=0. And again, split the equation into 2 parts: (x-1)=0 and (x-8)=0 And the two values of x you will get are 1 and 8